OpenStudy (anonymous):

Use the distributive property to simplify the expression below: −2(3d − 1) + 4d

OpenStudy (aroub):

why dont you use it? or you dont know?

OpenStudy (asnaseer):

distributive property means:\[a(b+c)=ab+ac\]you should be able to use this to simplify your expression.

OpenStudy (alfie):

a*(b+c) = a*b + a*c.